# Leopard[Uncaptioned image]: A Vision Language Model for Text-Rich Multi-Image Tasks

### v2

@@ Abstract
Text-rich images, where text serves as the central visual element guiding the overall understanding, are prevalent in real-world applications, such as presentation slides, scanned documents, and webpage snapshots. Tasks involving multiple text-rich images are especially challenging, as they require not only understanding the content of individual images but reasoning about inter-relationships and logical flows across multiple visual inputs. Despite the importance of these scenarios, current multimodal large language models (MLLMs) struggle to handle such tasks due to two key challenges: (1) the scarcity of high-quality instruction tuning datasets for text-rich multi-image scenarios, and (2) the difficulty in balancing image resolution with visual feature sequence length. To address these challenges, we propose Leopard, a MLLM designed specifically for handling vision-language tasks involving multiple text-rich images. First, we curated about one million high-quality multimodal instruction-tuning data, tailored to text-rich, multi-image scenarios. Second, we developed an adaptive high-resolution multi-image encoding module to dynamically optimize the allocation of visual sequence length based on the original aspect ratios and resolutions of the input images. Experiments across a wide range of benchmarks demonstrate our model’s superior capabilities in text-rich, multi-image evaluations and competitive performance in general domain evaluations. Our code is available at https://github.com/Jill0001/Leopard.
@@ 1 Introduction
Figure 1: Left: A demonstration of a text-rich multi-image task. Models need to reason about the textual content across multiple images to answer the question correctly. Leopard  successfully generates the right answer while baselines fail. Right: Evaluation results of Leopard and three baselines. Our model surpasses its counterparts across text-rich multi-image benchmarks by a large margin, maintaining comparable performance on single and general evaluations.
@@
Multimodal large language models (MLLMs) have revolutionized vision-language tasks, driving advancements in a variety of areas such as image captioning and object detection [67, 74, 72]. These improvements extend to applications involving text-rich images where text serves as the primary visual element guiding image comprehension, such as visual document understanding [51] and scene text recognition [59]. Traditional OCR-based pipelines in these text-rich visual scenarios are being replaced by end-to-end approaches that directly encode intertwined multimodal inputs [69, 75, 63], leading to improved efficiency and accuracy in handling text-rich images.
@@
Despite these advancements, the majority of existing open-source MLLMs, like LLaVAR [75] and mPlug-DocOwl-1.5 [16], have primarily focused on optimizing performance for text-rich single-image tasks. This focus inherently limits their applicability in many real-world scenarios, where tasks often involve multiple inter-connected images. For instance, multi-page visual document understanding requires integrating information spread across different pages to capture the logical flow across the whole document [64, 25]. To understand presentation slides, grasping the overarching narrative requires understanding multiple slides with unique but interrelated content [62]. These vision-language tasks on multiple text-rich images require advanced capabilities that go beyond merely recognizing text and visuals within a single image; they involve understanding and reasoning about relationships and logical flows across multiple visual inputs. While some models – such as OpenFlamingo [3], VILA [38], Idefics2 [28] – have made strides toward supporting multi-image inputs, they mainly focus on scenarios with natural images but fall short in understanding sequences of text-rich images with interrelated textual and visual information. We plot the performance of representatives of the aforementioned models in Figure 1. Upon examining their training data and model architecture, we identified two primary limitations.
@@
First, there is a scarcity of high-quality instruction tuning datasets on text-rich multi-image scenarios. Existing visual instruction tuning datasets for text-rich images are predominantly based on single-image inputs [22, 59, 48, 63], which limits the model ability to generalize and reason across multiple images. Second, in text-rich multi-image scenarios, there is a challenge of balancing image resolution and sequence length limitations. Many general-domain MLLMs adopt the low-resolution settings of pre-trained visual encoders [38, 21]. However, for text-rich images, such as scientific reports, recognizing text content becomes difficult at low resolutions. While some approaches overcome this in single-image settings by splitting the original image to preserve high-resolution details [40, 16], this approach is less effective when applied to multiple images, as it quickly exceeds model’s maximum sequence length. Moreover, compressing such long-sequence representations into shorter ones leads to significant information loss, thereby degrading model performance [3, 27]. Thus, a critical balance must be struck between maintaining sufficient visual detail and keeping sequence lengths manageable.
In this paper, we introduce a novel multimodal large language model, named Leopard [^1] Leopards have remarkable visual adaptations that allow them to track prey both from afar and up close, making them highly efficient hunters.. Leopard is specifically designed to handle complex text-rich, multi-image tasks. To train Leopard, we first curated about one million high-quality multimodal instruction-tuning data, tailored to the text-rich, multi-image scenarios. This dataset spans three key domains that are commonly encountered in real-world scenarios: (1) multi-page documents, (2) multi-charts and multi-tables, (3) webpage trajectories. These scenarios capture the increasing complexity and multimodal nature of modern digital information. In addition, to enable high-resolution encoding in multi-image inputs, we equipped Leopard with an adaptive high-resolution multi-image encoding module. Specifically, it dynamically optimizes the allocation of visual sequence length based on the original aspect ratios and resolutions of the input images. We then apply pixel shuffling to losslessly compress [8] long visual feature sequences into shorter ones. This approach allows the model to accommodate multiple high-resolution images without compromising detail or clarity.
@@
We conducted experiments on 13 vision-language benchmark datasets, evaluating Leopard from multiple perspectives. Consistent improvements were observed when training Leopard with two distinct base model architectures: LLaVA and Idefics2. Our results demonstrate Leopard’s superior performance on 5 text-rich, multi-image benchmarks, outperforming the best open-source MLLM by an average of +9.61 points. Moreover, Leopard remains highly competitive in text-rich single-image tasks and general-domain vision-language benchmarks, achieving comparable results to state-of-the-art MLLMs without extensive fine-tuning. Further ablation studies confirm the effectiveness of our instruction-tuning dataset and the adaptive high-resolution encoding module. These findings highlight Leopard’s strong performance across various multimodal applications.
@@ 2 Related Work
Multimodal Large Language Models (MLLMs). Many approaches have been proposed for building MLLMs, leveraging different architectural designs. A widely adopted approach is the decoder-only architecture, exemplified by LLaVA [41], Emu2 [61], and Intern-VL [9]. These models typically incorporated a visual encoder to encode images, a vision-language connector to project visual features into the language feature space, and a language model that processes both visual and textual information jointly. Another line of work employed cross-attention architectures where encoded image features are integrated with textual tokens via cross-attention layers, as seen in Flamingo [1], OpenFlamingo [3] and CogVLM [66]. Such a design allows models to retain the benefits of a fully intact language model but introduces new parameters to manage the visual-textual interplay.
Text-rich MLLMs. Text-rich images are traditionally processed in pipelines [60, 18], where an OCR module first recognized text from the image, followed by processing through a language model. To improve efficiency and avoid error propagation, with the advent of MLLMs, end-to-end approaches become more popular recently. For instance, LLaVAR [75] utilized a dataset of 400K instances with OCR-enhanced text to outperform LLaVA on various text-rich VQA tasks. Subsequent models such as UReader [70], TextMonkey [44], and Mplug-DocOwl-1.5 [16] recognized the importance of high-resolution encoding for accurate text comprehension, so they adopted strategies that cropped single images into multiple sub-images to preserve the original resolution during visual encoding. However, these approaches are primarily trained on single-image data, and struggle to generalize effectively to multi-image scenarios. Furthermore, the straightforward partitioning technique encounters challenges with multi-image inputs, as the sequence length rapidly increases with the number of images.
Multi-image MLLMs. Efforts have been made in training MLLMs with multi-image inputs due to the prevalence of multi-image scenarios in real-world applications. Mantis [21] introduced a multi-image instruction tuning dataset on a variety of natural image scenarios. Besides, both VILA [38] and Idefics-2 [28] incorporated image-text interleaved data during their pre-training. LLaVA-Next-Interleave [33] further extended this by incorporating videos and multi-view 3D data into the training pipeline. However, these works primarily target natural images and general visual understanding, leaving a gap in handling text-rich, multi-image scenarios. Natural images typically follow a different distribution from text-rich images and often do not demand high-resolution processing. As a result, many existing multi-image MLLMs struggle to generalize to text-rich scenarios. Our work aims to address this gap by specifically focusing on multi-image settings where text-rich images are the primary input. Very recently (in 08/2024 and 09/2024), multi-image training for MLLMs has attracted intense attention from researchers. Several concurrent efforts have included multi-image interleaved data to train their models, such as LLaVA-OneVision 08/2024 [31], Idefics3 (08/2024, 26), NVLM (09/2024, 10), mPlug-DocOwl-2 (09/2024, 17), Molmo (09/2024, 11) and Qwen2-VL (09/2024, 65). This trending paradigm highlights the significant practical value of multi-image MLLMs by enhancing their ability to tackle a wide range of real-world applications. The incorporation of multi-image instruction tuning data is therefore of paramount importance.
@@ 3 Method
Leopard follows the typical design of decoder-only vision language models [41, 40, 33], including a visual encoder, a vision language connector, and a language model (LM), as shown in Figure 2 (④⑤). Specially, the input images are first passed through the visual encoder, which extracts high-level visual features and captures essential semantic information. These visual features are then projected into the language representation space via the vision-language connector. After this transformation, the visual tokens are interleaved with the textual tokens, resulting in a sequence of interleaved text-visual tokens. This interleaved sequence is then fed into the LM, which processes these inputs in a causal manner, leveraging the contextual dependencies between text and visual information to generate coherent outputs that align with both modalities.
3.1 Multi-image text-rich Instruction Turning
Dataset
To train Leopard, we construct a large instruction-tuning dataset named Leopard-instruct, comprising 925K instances, with 739K specifically designed for text-rich, multi-image scenarios. While we extensively surveyed existing open-source datasets, we only identified 154K usable text-rich, multi-image samples, which is far from sufficient for effective instruction tuning, as shown in prior MLLM studies [21, 28, 33]. To address this data scarcity, we developed several data collection pipelines to collect high-quality text-rich, multi-image data, resulting in additional 585K instances. Each instance consists of a set of images along with corresponding task instructions and responses. The dataset details are presented in Table 1, and a detailed breakdown of its composition can be found in Appendix A.1.
Documents and Slides are common sources of multi-image data that primarily contain text and require cross-page context integration to fully understand the information.
These data is collected in three ways. First, we include 69K public multi-page document and slide datasets [64, 25, 79, 62], covering a variety of document types such as scanned handwriting, printed documents, and digital PDFs. Second, we adapt two single-page document datasets, DocVQA [51] and ArxivQA [34], for multi-image settings. Following [21], we randomly merge 2 to 4 single-page instances by concatenating their respective images and Q-A pairs. Prompts like “in the second image” are added to direct the model’s focus to the appropriate image. These merged samples help the model learn how natural language references align with corresponding image features. Third, we collect raw slides from [57] and SlideShare [^2] https://www.slideshare.net, and use GPT-4o to generate Q-A pairs and reasoning steps. We show the prompt to GPT in Figure 5. Upon manually reviewing 100 instances annotated by GPT-4o, we found an accuracy rate over 90%, indicating high annotation quality.
Tables and Charts provide highly organized, structured quantitative information, often involving complex data patterns and relationships, requiring the integration of both visual and textual elements for accurate interpretation.
To address the lack of instruction tuning data involving multiple tables or charts, we use the following strategies. First, we include 21K open-source multi-chart and multi-table datasets [77, 54], originally stored in JSON or DataFrame formats. We programmatically render these tables as images, converting them into multimodal data. Details of rendering can be found in Appendix A.3 Second, We utilize the TableGPT [35] dataset and split each table into multiple sub-tables, then convert them into figures, thereby creating multi-modal, multi-table instruction data. Third, we apply the same merging strategy used for combining single-page documents to synthesize multi-image datasets. This approach integrates several single-chart datasets, including ChartGemma [49], ChartQA [48], DVQA [22], and FigureQA [23]. Besides, we generate new multi-chart data from social reports of the Pew Research Center [^3] https://www.pewresearch.org that feature multiple interrelated charts within the articles under the same topic. We download charts from the website and use GPT-4o to create 20K Q-A pairs that require multi-chart understanding.
@@ 3.1 Multi-image text-rich Instruction Turning
Webpage Snapshots consist of sequential images representing web pages, providing visual context for user interactions and task flows. Understanding webpage is a critical skill for MLLMs to evolve into fully autonomous web agents [12, 14]. To collect relevant data, we format several web-related multimodal datasets into a Q-A structure as follows:
  1. 1.
    Web action prediction data: We include Mind2Web [12] and OmniACT [24], where we divide long web snapshots into multiple sub-figures, and plot bounding boxes based on the coordinates of web elements. Then GPT-4o is used to convert the original action data into a Q-A format, where the task is to identify the correct element to interact with.
  2. 2.
    Web-based classification data: We incorporate WebScreenshots [4], WebVision [36], and WebUI [68]. We utilize the web snapshots in these datasets and employ GPT-4o to generate Q-A pairs on webpage understanding, including chain-of-thought reasoning steps. The prompting details are provided in Figure 6.
@@
Otter-9B [30] CLIP ViT-L $224^{2}$ LLaMA-7B 9B 30M 5.1M
Emu2-Chat [61] EVA-02-CLIP $448^{2}$ LLaMA-33B 37B - 160M
MM1-7B-Chat [52] CLIP ViT-H $378^{2}$ 7B - 1.5M
VILA1.5-8B [38] SigLIP $384^{2}$ LLaMA3-8B 8B 50M 1M
mPlug-DocOwl-1.5 [16] CLIP ViT-L $448^{2}$ (x9 crops) LLaMA-7B 8B 4M 1M
Idefics2-8B [28] SigLIP $980^{2}$ Mistral-7B 8B 350M 20M
LLaVA-NeXT-Inter [33] SigLIP AnyRes Qwen1.5-7B 7B 1.3M 1.2M
Mantis-LLaVA [21] SigLIP $384^{2}$ LLaMA3-8B 8B 0.5M 1M
Mantis-Idefics2 [21] SigLIP $980^{2}$ Mistral-7B 8B 350M 1M
@@
Table 2: A detailed comparison of the model training details, including image resolution, vision encoder, backbone LLM, number of parameters (Param.), pre-training (PT.) data size, and instruction tuning (IT.) data size of baselines. AnyRes denotes the resolution selecting method proposed by [40] and Adapt HR. represents the proposed adaptive high-resolution multi-image encoding strategy.
@@
Augmenting with Rationales. In contrast to single-image tasks, multi-image scenarios typically require MLLMs to integrate information across multiple images, making cross-image reasoning difficult to train when only the final answer is provided [78, 19]. To address this, we employ GPT-4o to generate chain-of-thought (CoT) rationales for inherently multi-image datasets (excluding those formed by merging single-image data) that lack CoT annotations. This results in 250K instances with GPT-annotated reasoning, with the prompt detailed in Figure 7.
Other Domains. We also include datasets from various other domains such as maps (MapQA, 6), infographics (InfographicVQA, 50), mathematical diagrams (MathV360K, 58), and abstractive diagrams (IconQA, 47). We also incorporate mixed-domain datasets for text-rich images, including LLaVAR [75], Monkey, 37, and mPlugDocReason [16]. We remove duplicate subsets from these mixed-domain datasets. Among these datasets, 64K samples consist of multi-image data, while the remaining are single-image samples. To preserve natural image understanding ability, we add 313K samples from ShareGPT4V [7], an instruction dataset for natural images.
3.2 Adaptive High-resolution Multi-image
Encoding
Image resolution significantly influences the visual perception and understanding capabilities of MLLMs, particularly when processing text-rich images. Low-resolution images often cause printed text to become blurred or unreadable, resulting in misinterpretations, perception errors, and visual hallucinations. The visual resolution of most existing MLLMs is determined by their pre-trained visual encoders, which are typically limited to low resolutions such as $224\times 224$ or $336\times 336$ pixels [39, 38, 21]. These low-resolution constraints can hinder MLLMs to accurately understand textual information embedded within images.

### v3

# Leopard: A Vision Language Model for Text-Rich Multi-Image Tasks

@@ Abstract
Text-rich images, where text serves as the central visual element guiding the overall understanding, are prevalent in real-world applications, such as presentation slides, scanned documents, and webpage snapshots. Tasks involving multiple text-rich images are especially challenging, as they require not only understanding the content of individual images but reasoning about inter-relationships and logical flows across multiple visual inputs. Despite the importance of these scenarios, current multimodal large language models (MLLMs) struggle to handle such tasks due to two key challenges: (1) the scarcity of high-quality instruction tuning datasets for text-rich multi-image scenarios, and (2) the difficulty in balancing image resolution with visual feature sequence length. To address these challenges, we propose Leopard, an MLLM tailored for handling vision-language tasks involving multiple text-rich images. First, we curated about one million high-quality multimodal instruction-tuning data, tailored to text-rich, multi-image scenarios. Second, we proposed an adaptive high-resolution multi-image encoding module to dynamically optimize the allocation of visual sequence length based on the original aspect ratios and resolutions of images.
Experiments on a diverse set of benchmarks reveal that our model consistently outperforms state-of-the-art systems, such as Llama-3.2 and Qwen2-VL, in challenging text-rich, multi-image evaluations. Remarkably, our approach achieves outstanding performance using only 1.2M training instances, all of which are fully open-sourced, demonstrating both high efficiency and effectiveness compared to models trained on large-scale in-house data. Our code and data are available at https://github.com/tencent-ailab/Leopard.
@@ 1 Introduction
Multimodal large language models (MLLMs) have revolutionized vision-language tasks, driving advancements in various areas such as image captioning and object detection (Zhang et al., 2024; Zang et al., 2024). These improvements extend to applications involving text-rich images where text serves as the primary visual element guiding image comprehension, such as visual document understanding (Mathew et al., 2021) and scene text recognition (Singh et al., 2019b). Traditional OCR-based pipelines in these text-rich visual scenarios are being replaced by end-to-end approaches that directly encode intertwined multimodal inputs (Wu et al., 2023b; Zhang et al., 2023; Tang et al., 2024), leading to improved accuracy in handling text-rich images.
Figure 1: Left: A demonstration of a text-rich multi-image task, where models must reason across multiple images to answer correctly. Leopard generates the correct answer, while baselines fail. Right: Leopard outperforms three baselines on text-rich multi-image benchmarks by a large margin, while maintaining comparable performance on single-image and general evaluations.
@@
Despite these advancements, the majority of existing open-source MLLMs, like LLaVAR (Zhang et al., 2023) and mPlug-DocOwl-1.5 (Hu et al., 2024a), have primarily focused on optimizing performance for text-rich single-image tasks. This focus inherently limits their applicability in many real-world scenarios, where tasks often involve multiple inter-connected images. For instance, multi-page visual document understanding requires integrating information spread across different pages to capture the logical flow across the whole document (Tito et al., 2022; Landeghem et al., 2023). To understand presentation slides, grasping the overarching narrative requires understanding multiple slides with unique but interrelated content (Tanaka et al., 2023). These vision-language tasks on multiple text-rich images require advanced capabilities that go beyond merely recognizing text and visuals within a single image; they involve understanding and reasoning about relationships and logical flows across multiple visual inputs. While some models – such as OpenFlamingo (Awadalla et al., 2023), VILA (Lin et al., 2023), Idefics2 (Laurençon et al., 2024b) – have made strides toward supporting multi-image inputs, they mainly focus on scenarios with natural images but fall short in understanding sequences of text-rich images with interrelated textual and visual information. We plot the performance of representatives of the aforementioned models in Figure 1. Upon examining their training data and model architecture, we identified two primary limitations.
@@
First, there is a scarcity of high-quality instruction tuning datasets on text-rich multi-image scenarios. Existing visual instruction tuning datasets for text-rich images are predominantly based on single-image inputs (Kafle et al., 2018; Singh et al., 2019b; Masry et al., 2022; Tang et al., 2024), which limits the model ability to generalize and reason across multiple images. Second, in text-rich multi-image scenarios, there is a challenge of balancing image resolution and sequence length limitations. Many general-domain MLLMs adopt the low-resolution settings of pre-trained visual encoders (Lin et al., 2023; Jiang et al., 2024). However, for text-rich images, such as scientific reports, recognizing text content becomes difficult at low resolutions. While some approaches overcome this in single-image settings by splitting the original image to preserve high-resolution details (Liu et al., 2024a; Hu et al., 2024a), this approach is less effective when applied to multiple images, as it quickly exceeds model’s maximum sequence length. Moreover, compressing such long-sequence representations into shorter ones leads to significant information loss, thereby degrading model performance (Awadalla et al., 2023; Laurençon et al., 2023). Thus, a critical balance must be struck between maintaining sufficient visual detail and keeping sequence lengths manageable. In this paper, we introduce a novel multimodal large language model, Leopard, designed for complex text-rich, multi-image tasks. To train Leopard, we curated approximately one million high-quality multimodal instruction-tuning samples, tailored for text-rich, multi-image contexts. This dataset was developed through a combination of new multi-image data collection from the web with augmentation of question-answer pairs, transformation of unimodal text data into multi-image formats, and reformatting single-image resources by employing an assembly strategy. Ultimately, the dataset encompasses three essential domains prevalent in real-world scenarios: (1) multi-page documents, (2) multi-charts and multi-tables, and (3) webpage sequences, reflecting the complex, multimodal demands of modern digital information. In addition, to enable high-resolution encoding in multi-image inputs, we equipped Leopard with an adaptive high-resolution multi-image encoding module. Specifically, it dynamically optimizes the allocation of visual sequence length based on the original aspect ratios and resolutions of the input images. We then apply pixel shuffling to losslessly compress (Chen et al., 2024) long visual feature sequences into shorter ones. This approach allows the model to accommodate multiple high-resolution images without compromising detail or clarity.
@@
We conducted experiments on 12 vision-language benchmark datasets, evaluating Leopard from multiple perspectives. Consistent improvements were observed when training Leopard with two distinct base model architectures: LLaVA and Idefics2. Our results demonstrate Leopard’s superior performance on 5 text-rich, multi-image benchmarks, outperforming the best open-source MLLM by an average of +8.5 points. Moreover, Leopard remains highly competitive in text-rich single-image tasks and general-domain vision-language benchmarks, achieving comparable results to state-of-the-art MLLMs without extensive fine-tuning. Further ablation studies confirm the effectiveness of our instruction-tuning dataset and the adaptive high-resolution encoding module. These findings highlight Leopard’s strong performance across various multimodal applications.
@@ 2 Related Work
Text-rich MLLMs. Text-rich images are traditionally processed in pipelines (Singh et al., 2019a; Hu et al., 2020), where an OCR module first recognized text from the image, followed by processing through a language model. To improve efficiency and avoid error propagation, with the advent of MLLMs, end-to-end approaches become more popular recently. For instance, LLaVAR (Zhang et al., 2023) utilized a dataset of 400K instances with OCR-enhanced text to outperform LLaVA on various text-rich VQA tasks. Subsequent models such as UReader (Ye et al., 2023), TextMonkey (Liu et al., 2024c), and Mplug-DocOwl-1.5 (Hu et al., 2024a) recognized the importance of high-resolution encoding for accurate text comprehension, so they adopted strategies that cropped single images into multiple sub-images to preserve the original resolution during visual encoding. However, these approaches are primarily trained on single-image data, and struggle to generalize effectively to multi-image scenarios. Furthermore, the straightforward partitioning technique encounters challenges with multi-image inputs, as the sequence length rapidly increases with the number of images.
Multi-image MLLMs. Efforts have been made in training MLLMs with multi-image inputs due to the prevalence of multi-image scenarios in real-world applications. Mantis (Jiang et al., 2024) introduced a multi-image instruction tuning dataset on a variety of natural image scenarios. Besides, both VILA (Lin et al., 2023) and Idefics-2 (Laurençon et al., 2024b) incorporated image-text interleaved data during their pre-training. LLaVA-Next-Interleave (Li et al., 2024c) further extended this by incorporating videos and multi-view 3D data into the training pipeline. However, these works primarily target natural images and general visual understanding, leaving a gap in handling text-rich, multi-image scenarios. Natural images typically follow a different distribution from text-rich images and often do not demand high-resolution processing. As a result, many existing multi-image MLLMs struggle to generalize to text-rich scenarios. Our work aims to address this gap by specifically focusing on multi-image settings where text-rich images are the primary input. Very recently, multi-image training for MLLMs has attracted intense attention from researchers. Several concurrent efforts have included multi-image interleaved data to train their models, such as LLaVA-OneVision 08/2024 (Li et al., 2024b), Idefics3 (08/2024, (Laurençon et al., 2024a)), NVLM (09/2024, (Dai et al., 2024)), mPlug-DocOwl-2 (09/2024, (Hu et al., 2024b)), Molmo (09/2024, (Deitke et al., 2024)) and Qwen2-VL (09/2024, (Wang et al., 2024)). This trending paradigm highlights the significant practical value of multi-image MLLMs by enhancing their ability to tackle a wide range of real-world applications. The incorporation of multi-image instruction tuning data is therefore of paramount importance.
@@ 3 Method
Leopard follows the typical design of decoder-only vision language models (Liu et al., 2023b; 2024a; Li et al., 2024c), including a visual encoder, a vision language connector, and a language model (LM), as shown in Figure 2 (④⑤).
@@
Specially, the input images are first passed through a visual encoder to extract high-level semantic features, which are then mapped into the language space via a vision-language connector. After this transformation, the visual tokens are interleaved with the textual tokens and fed into the language model, which processes them causally, leveraging text-visual context to generate coherent outputs that align with both modalities.
3.1 Multi-image Text-rich Instruction Dataset
To train Leopard, we create a large-scale instruction-tuning dataset named Leopard-instruct, comprising 925K instances, with 739K specifically tailored for text-rich, multi-image scenarios. Table 1 lists the composition of our data, with a detailed breakdown in the Appendix A.1.
In constructing Leopard-instruct, we surveyed existing open-source datasets but found only 154K suitable instances containing text-rich, multi-image data. This volume was insufficient for effective instruction tuning, compared to prior MLLM studies (Jiang et al., 2024; Laurençon et al., 2024b; Li et al., 2024c). To address this data scarcity, we collect an additional 585K high-quality instances of text-rich, multi-image data. Each instance consists of a set of images paired with task instructions and responses. To collect these instances, we design multiple data pipelines, such as collecting multi-image data from the web, transforming textual data into multi-image data, and assembling multi-turn multi-image data. These methods are discussed in detail in the following paragraphs.
3.1.1 Newly Collected Multi-image Data (585K)
Web Data Annotation: We curate data from websites containing text-rich, multi-image content, such as articles featuring multiple charts or presentation slide decks. Multiple charts from the same article are often semantically interconnected, requiring sophisticated cross-image reasoning for comprehension. Currently, no datasets are specifically designed for multi-chart scenarios. To address this, we download charts from social reports of the Pew Research Center [^1] https://www.pewresearch.org which typically feature multiple, interrelated charts on the same topic. Following a common approach (Shi et al., 2024), we use GPT-4o to generate questions about these charts, as well as the corresponding answers. To further enhance Leopard’s reasoning abilities, we prompt GPT-4o to create questions that necessitate cross-chart information synthesis and to provide step-by-step reasoning explanations before reaching the final answer. Slide decks offer another valuable source of text-rich, multi-image data, as combining information across multiple slides are often needed for complete understanding. We download slides from SlideShare [^2] https://www.slideshare.net, and again use GPT-4o to generate question-answer pairs, along with detailed reasoning steps. The GPT-4o annotation prompt is provided in the Appendix A.3. After manually reviewing a sample of 100 GPT-4o-annotated instances, we observed an accuracy rate exceeding 90%, indicating high-quality annotations.
@@ 3.1.1 Newly Collected Multi-image Data (585K)
Otter-9B (Li et al., 2023) CLIP ViT-L $224^{2}$ LLaMA-7B 9B 30M 5.1M
Emu2-Chat (Sun et al., 2023) EVA-02-CLIP $448^{2}$ LLaMA-33B 37B - 160M
MM1-7B-Chat (McKinzie et al., 2024) CLIP ViT-H $378^{2}$ 7B - 1.5M
VILA1.5-8B (Lin et al., 2023) SigLIP $384^{2}$ LLaMA3-8B 8B 50M 1M
mPlug-DocOwl-1.5 (Hu et al., 2024a) CLIP ViT-L $448^{2}$ (x9 crops) LLaMA-7B 8B 4M 1M
Idefics2-8B (Laurençon et al., 2024b) SigLIP $980^{2}$ Mistral-7B 8B 350M 20M
LLaVA-NeXT-Inter (Li et al., 2024c) SigLIP AnyRes Qwen1.5-7B 7B 1.3M 1.2M
Llama-3.2-11B (Meta et al., 2024) ViT-H/14 $384^{2}$(x4 crops) LLaMA3.1-8B 11B 6B -
Qwen2-VL-7B (Wang et al., 2024) CLIP ViT-bigG Dynamic Qwen2-7B 7B 800B Token -
Mantis-LLaVA (Jiang et al., 2024) SigLIP $384^{2}$ LLaMA3-8B 8B 0.5M 1M
Mantis-Idefics2 (Jiang et al., 2024) SigLIP $980^{2}$ Mistral-7B 8B 350M 1M
@@
Leopard-LLaVA-Pro (Ours) SigLIP Adapt HR. LLaMA3.1-8B 8B 3.6M 1.2M
@@
Table 2: A detailed comparison of the model training details between baseline models and Leopard, including image resolution, vision encoder, backbone LLM, number of parameters (Param.), pre-training (PT.) data size, and instruction tuning (IT.) data size of baselines. AnyRes denotes the resolution selecting method proposed by Liu et al. (2024a) and Adapt HR. represents the proposed adaptive high-resolution multi-image encoding strategy. We show the instance numbers (or token number if reported) for training of the models.
Transformed from Unimodal Textual Data: Tables provide highly structured, organized quantitative information that requires both visual layout perception and textual understanding. However, existing vision-language tasks rarely address multi-table understanding. To fill this gap, we employ two strategies. First, we incorporate multi-table data from MultiHiertt (Zhao et al., 2022) and MultiTabQA (Pal et al., 2023) where tables are originally stored in JSON or DataFrame formats. We programmatically render tables as images to convert them into multimodal data. Details of the rendering process are in Appendix A.4. Second, we use single-table data from TableGPT (Li et al., 2024e) and split each table into sub-tables by randomly dividing rows and columns. These sub-tables are then rendered as individual images, creating multimodal, multi-table instruction data.
Assembled Multi-turn Multi-image Data: Single-image datasets are significantly more abundant than multi-image resources. To leverage such data in training Leopard, inspired by Jiang et al. (2024), we randomly combine 2 to 4 single-image instances to create synthetic multi-image data. We first concatenate the selected images to form a multi-image input, then stack their corresponding question-answer pairs as multi-turn conversations. Prompts like “in the second image” and “from the image on the right-hand”, are added to each turn to direct the model’s focus to the appropriate image. These assembled samples enhance the model’s ability to associate natural language references with corresponding visual features but also expand the model’s exposure to diverse domains beyond the limited multi-image datasets. This assembly strategy is applied on single-image datasets including ArxivQA Li et al. (2024d), RICO Deka et al. (2017), FigureQA Kahou et al. (2018), and MapQA Chang et al. (2022).
Dataset Total #Text Domain Anno. Summary
Interleave (M4-Instruct) (Li et al., 2024c) 500.8K 21.3K Natural images (real-world scenes, animals, landscapes, people, etc.) N/A Contains limited text-rich images; primarily focused on natural scenes.
Mantis-Instruct (Jiang et al., 2024) 721K 0 Natural images (real-world scenes, animals, landscapes, people, etc.) GPT-4V No text-rich images included.
MP-DocStruct1M (Hu et al., 2024b) 1.113M 1.113M Documents only N/A Contains only document images; single domain with low diversity.
MIMIC-IT (Li et al., 2023) 2.8M 0 Natural images (real-world scenes, animals, landscapes, people, etc.) N/A No text-rich images included.
Leopard-Instruct 739K 739K Text-rich images (documents, charts, tables, webpages, slide decks, etc.) GPT-4o Diverse text-rich images across various domains, with silver labels by GPT-4o.
Table 3: Comparison of 5 multi-image instruction datasets. Total denotes the total number of multi-image samples and # Text is the text-rich samples number. Anno. denotes if the dataset has annotations.
@@
Augmenting with Rationales. Unlike single-image tasks, multi-image scenarios often require MLLMs to aggregate information across multiple images. However, many existing datasets provide only final answers (Zheng et al., 2023; Hu et al., 2023), limiting the model’s ability to learn cross-image reasoning skills. To address this limitation, we employ GPT-4o to generate chain-of-thought (CoT) rationales for multi-image datasets that lack such annotations. This results in 250K instances with GPT-annotated reasoning steps. (details in Appendix A.3)
3.1.2 Adopted Existing Data (340K)
We comprehensively investigate existing text-rich training resources and adopt the following data, including both multi-image and single-image instances from various domains. Documents are a common source of multi-image data, containing extensive text that requires cross-page context integration to fully capture the intended information. We include public multi-page document datasets (Tito et al., 2022; Landeghem et al., 2023; Zhu et al., 2022), covering a variety of document types such as scanned handwriting, printed documents, and digital PDFs. Slide presentations, similarly, include text-rich content spread across multiple pages. Besides downloading new data from the web, we use existing slides datasets Tanaka et al. (2023); Sefid et al. (2021). Webpage snapshots are composed of sequential images of webpages and provide visual context essential for interpreting user instructions on web-based tasks. We employ both web action prediction data (Mind2Web (Deng et al., 2023), OmniACT (Kapoor et al., 2024)) and web-based question answering data (WebScreenshots (Aydos, 2020), WebVision (Li et al., 2017), WebUI (Wu et al., 2023a)). Data in other domains are also included, such as infographics data (InfographicVQA Mathew et al. (2022)), mathematical diagrams (MathV360K Shi et al. (2024)), and abstractive diagrams (IconQA Lu et al. (2021)). We also incorporate some mixed-domain datasets that contain different types of text-rich images, including LLaVAR (Zhang et al., 2023), Monkey Li et al. (2024g), and mPlugDocReason (Hu et al., 2024a). To preserve natural image understanding ability, we add samples from an instruction tuning dataset for natural images – ShareGPT4V (Chen et al., 2023). Finally, we remove duplicate instances from the whole Leopard-instruct dataset.
3.1.3 Comparison with Existing Datasets
We comprehensively compare several widely-used multi-image instruction-tuning datasets with Leopard-instruct in Table 3. Existing datasets such as Interleave (M4-Instruct), despite containing text-rich images, offer a very limited quantity (21.3K samples), insufficient for extensive training. While MP-DocStruct1M presents a substantial volume of document-focused data (1.113M samples), its scope remains narrow, limited exclusively to documents. Other datasets like Mantis-Instruct and MIMIC-IT primarily consist of natural images, limiting their applicability in text-rich scenarios. Addressing these limitations, our proposed Leopard-instruct dataset provides a robust collection of 739K text-rich images spanning diverse domains—including documents, charts, tables, webpages, and slide decks. This diversity and scale make Leopard-instruct particularly useful for training models to handle text-rich multi-image tasks.
3.2 Adaptive Resolution Multi-Image Encoding
Image resolution plays a critical role in MLLMs’ visual comprehension, especially for text-rich images. Low resolutions can blur printed text, leading to perception errors and visual hallucinations. Existing MLLMs rely on pre-trained visual encoders typically capped at low resolutions like $224\times 224$ or $336\times 336$ pixels (Liu et al., 2023a; Lin et al., 2023; Jiang et al., 2024), limiting accurate textual understanding within images.